OpenStudy (anonymous):

Evaluate |-6| + 3|4| + |-5| - 0 · 6. 0 18 23 25

OpenStudy (anonymous):

Like I said in the last question. Absolute value turns negative numbers into positive and does not affect positive numbers at all

OpenStudy (anonymous):

so it turns into 6+3(4)+5-0(6)=

OpenStudy (anonymous):

first off. Do you know what PEMDAS is stands for?

OpenStudy (anonymous):

@GirlByte nope

OpenStudy (anonymous):

It is used to properly do multiple problems. Parenthesis Exponents Multiply Divide Add Subtract

OpenStudy (anonymous):

You have to do these in order to get the right answer

OpenStudy (anonymous):

So do you know what to do first?

OpenStudy (anonymous):

Do you understand what im saying? tell me if you are confused ^_^

OpenStudy (anonymous):

Dont be shy if you get it wrong xP

OpenStudy (anonymous):

you don whats in the parenthesis first then exponent then either multiply, divide or add right ???

OpenStudy (anonymous):

@GirlByte

OpenStudy (anonymous):

Yes. But in this case we do not have parenthesis with two numbers in it and exponents so we skip those steps

OpenStudy (anonymous):

So what do we do to this problem? 6+3(4)+5-0(6)=

OpenStudy (anonymous):

Remember: If two numbers are beside each other like this 2(3). you need to multiply them

OpenStudy (anonymous):

okay okay yeah i remeber that now lol i had a brain fart

OpenStudy (anonymous):

Good lol

OpenStudy (anonymous):

Ok since we dont have parenthesis and exponents WE need to ____

OpenStudy (anonymous):

Multiply

OpenStudy (anonymous):

6+3(4)+5-0(6)= 6+12+5-0=

OpenStudy (anonymous):

then add

OpenStudy (anonymous):

the answer is 23